Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hour and train B is traveling at 80 miles per hour. Train A passes a station at 3:25 am. If train B passes the same station at 3:40 am, at what time will train B catch up with train A?
Answer by mananth(16949)

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60 miles per hour and train B is traveling at 80 miles per hour. Train A passes a station at 3:25 am. If train B passes the same station at 3:40 am, at what time will train B catch up with train A?
...
Train A speed 60 mph
train B 80 mph.
Train A passes a station at 3:25 am.
If train B passes the same station at 3:40 am,
..
Speed train B - speed train A = catch up speed
80-60 = 20 mph.
..
time between trains passing the station 3.40 -3.25=15 minutes = 1/4 hour
..
in 1/4 hour train A travesl 60*1/4 = 15 miles.
...
catch up time = catch up distance / catch up speed
catch up time = 15/20= 3/4 hours
..
Train B will catch up with Train A at 3.40 + 0.45= 4.25 pm
..
m.ananth@hotmail.ca